OpenStudy (anonymous):

The length of a rectangle is five times its width. If the area of the rectangle is 180m^2, find its perimeter.

OpenStudy (anonymous):

You are given with; L = 5W So: Area is given by: \[Area = L \times W\] \[180 = L \times W\] \[180 = 5 W^2 \implies W^2 = 36\] Find W here.. Then to find perimeter use: \[Perimeter = 2(L + W)\]

OpenStudy (anonymous):

x(5x)=180 5x2=180 x=36 2x+2(5x)=72+360=432 ANSWER = 36

OpenStudy (anonymous):

answer is not 36.

OpenStudy (anonymous):

a*5a =180m^2 a= square (36) m a = 6 perimeter = 5*6 +5 = 36 meters, my guess

OpenStudy (anonymous):

Multiplied by 2 instead..

OpenStudy (anonymous):

hmmm okay ..... ill redo and see if i made a mistake

OpenStudy (anonymous):

yeah you are right 72 is the answer since it's 2 (with + length)

OpenStudy (anonymous):

W = 6 here.. So L = 30 here.. Just add them and then Multiply by 2, to get Perimeter..